# Testing New Physics in Oscillations at a Neutrino Factory

### v1

@@ 3.2 Results for non-standard vector interactions
For a more general case, we take the off-diagonal NSI parameters to be complex and marginalize over the corresponding phases in Fig. 3. Due to the free complex phase, the degeneracies can be enlarged and significantly reduce the experimental sensitivities to the NSI parameters. Especially, the degeneracy between $\epsilon_{ee}-\epsilon_{\mu\mu}$ and $|\epsilon_{e\tau}|$ leads to a wing-like structure for the original DUNE configuration which comes from the non-trivial dependence among the $\phi_{e\tau}$, $\delta_{CP}$ and $\theta_{23}$ parameters Coloma:2015kiu . To better understand this degeneracy we show in the left panel of Fig. 4 the oscillation probabilities without NSI and with non-zero $\epsilon_{ee}-\epsilon_{\mu\mu}$ and $\epsilon_{e\tau}$. For the standard case without NSI, we use the values of oscillation parameters from Tab. 1, while we take the best fit from that minimizes the $\chi^{2}$ function for the NSI case. More explicitly, we choose the point ($\epsilon_{ee}-\epsilon_{\mu\mu}=1.2$ and $|\epsilon_{e\tau}|=0.3$) in the wing-like structure panel of Fig. 3 as the input and a parameter set $\phi_{e\tau}=-152^{\circ}$, $\theta_{23}=42^{\circ}$, and $\delta_{CP}=-53^{\circ}$ the values that lead to a $\chi^{2}$ minimum, which indicates this parameter set has a degeneracy with the standard case without NSI. As DUNE only measures $\nu_{e}$ appearance and $\nu_{\mu}$ disappearance channels, this degeneracy significantly reduces DUNE’s sensitivity to these NSI parameters. It is exciting to see the addition of a neutrino factory can break this degeneracy and significantly enhance the sensitivity in the $\epsilon_{ee}-\epsilon_{\mu\mu}$-$|\epsilon_{e\tau}|$ parameter space. This is because a neutrino factory can measure the $\nu_{e}$ disappearance channel and the $\nu_{\mu}$ appearance channel, which help break the existing degeneracy in the $\nu_{e}$ appearance and $\nu_{\mu}$ disappearance channels.
@@
Figure 3: The experimental constraint on different combinations of two NSI parameters at 90% confidence level, where the corresponding complex phase of NSI off-diagonal parameter and $\delta_{\rm CP}$ are marginalized over.
@@
In addition to the expected constraints on NSI parameters, we also study the precision of the Dirac CP phase as well as the sensitivity to $\theta_{23}$ in the presence of NSI. To quantify the expected precision on to $\delta_{CP}$, we define $\Delta\delta_{CP}\equiv\rm max\left\{|\delta_{CP}^{\rm fit}-\delta_{CP}^{\rm true}|\right\}$ to be the maximum absolute value of the difference between $\delta_{CP}^{\rm fit}$ and $\delta_{CP}^{\rm true}$, where $\delta_{CP}^{\rm fit}$ is the fitted Dirac CP phase which gives $\Delta\chi^{2}=2.71$ (90% C.L.) and $\delta_{CP}^{\rm true}$ is the fixed true Dirac CP phase. We take the benchmark value of $\delta^{\rm true}_{CP}=0^{\circ}$ or $-90^{\circ}$, and three benchmark values of $\theta^{\rm true}_{23}=42^{\circ}$, $45^{\circ}$, or $49^{\circ}$ for completeness. The results for $\delta^{\rm{true}}_{CP}=0$ are shown in the appendix which show qualitatively similar improvements to the precision of $\delta_{CP}$ with the addition of a neutrino factory. Similar to the standard oscillation framework without NSI, the precision on $\delta_{CP}^{\rm true}=0^{\circ}$ is better than for the $\delta_{CP}^{\rm true}=-90^{\circ}$ case while the octant of $\theta_{23}$ has only a mild impact on the CP precision. In the $\delta_{CP}^{\rm true}=0^{\circ}$ case, the increase in precision by combing DUNE with a neutrino factory can be enhanced by a factor of 2 (1.5) compared with the original DUNE configuration in the presence of diagonal (off-diagonal) NSI parameter. A sizable improvement exists for $\delta_{CP}^{\rm true}=-90^{\circ}$, as seen in Fig. 5.
@@
In Fig. 6 we show the sensitivity to $\theta_{23}$ for different benchmark values of $\delta^{\rm true}_{CP}$ and $\theta^{\rm true}_{23}$ as discussed above. Here we take the $\epsilon_{e\mu}$ parameter for illustration since the neutrino factory has a significant improvement in this case, while the results for other NSI parameters are shown in the appendix. In the figure, $\epsilon_{e\mu}$ is marginalized over its modulus together with its phase while the true value is taken to be zero. For benchmark values of $\theta_{23}^{\rm true}=42^{\circ}$ and $49^{\circ}$, the $\theta_{23}$ sensitivity of DUNE can be significantly reduced if one also considers NSI. As a result, the $\theta_{23}$ octant cannot be determined with over $\sim 3\sigma$ C.L. at the benchmarks considered. With the addition of a neutrino factory, however, the sensitivity to $\theta_{23}$ octant even when minimizing over $\epsilon_{e\mu}$ with a phase can reach over $5\sigma$. We also find that the true value of $\delta_{CP}^{\text{true}}$ has only a small impact on the sensitivity.
@@
Figure 5: The expected uncertainty on $\delta_{CP}$ at 90% confidence level, where the corresponding NSI parameter is minimized. In this figure, the off-diagonal NSI parameters are taken as complex and the corresponding phase is minimized together with its modulus at a time. We assume $\delta_{CP}^{\text{true}}=-90^{\circ}$ and $\theta_{23}^{\text{true}}=42^{\circ},~{}45^{\circ},~{}49^{\circ}$. See Fig. 11 in the appendix for $\delta_{CP}=0$.
@@
Figure 6: The experimental sensitivity to $\theta_{23}$ for different true values of $\theta_{23}$ and we assume no NSI effect in the true case. For the fitting parameter, we consider the complex off-diagonal NSI parameter $\epsilon_{e\mu}$, which is minimized over including its complex phase. See Figs. 12-15 in the appendix.
@@ 4.1 Framework
The CPT symmetry, which connects three discrete symmetries: charge conjugation (C), parity (P), and time reversal (T), is one of the most essential symmetries in particle physics and its invariance has been a guiding tool to build models. The CPT theorem is based on several basic assumptions on models, including Lorentz invariance, hermiticity of the Hamiltonian, and local commutativity. CPT violation can appear in several classes of more complete models and may affect neutrino oscillations Kostelecky:1988zi ; Kostelecky:1989jp ; Kostelecky:1989jw ; Sheikh-Jabbari:2000hdt ; Carroll:2001ws ; Greenberg:2002uu ; Barenboim:2002tz ; DeGouvea:2002xp ; Ge:2019tdi . Certainly, the discovery of the CPT violation would indicate that at least one basic assumption should be rejected and all our model building strategy need to be revisited. Therefore, testing the CPT symmetry is one of the major tasks in particle physics.
@@
Here $x$ ($\bar{x}$) corresponds to the oscillation parameter set in the neutrino (antineutrino) oscillation probabilities.
Figure 7: The experimental sensitivity to the difference between the oscillation parameters for neutrino and anti neutrinos $x$-$\bar{x}$ at 90% confidence level for different experimental setups assuming CPT conservation. We mark the best-fit value with a star. The dashed line shows the CPT conserving line.
Figure 8: The expected experimental constraint on $\theta_{23}$-$\bar{\theta}_{23}$ at 90% confidence level. Left panel: both $\theta_{13}$ and $\bar{\theta}_{13}$ are free parameters without any prior. Right panel: both $\theta_{13}$ and $\bar{\theta}_{13}$ are free parameters, but only the prior of $\bar{\theta}_{13}$ is included.
@@
Figure 10: CPT violation benchmark plot. Upper left panel: $\sin^{2}\theta_{13}^{\rm true}=0.032,\sin^{2}\bar{\theta}_{13}^{\rm true}=0.022$. Upper right panel: $\theta_{23}^{\rm true}=49^{\circ},\bar{\theta}_{23}^{\rm true}=42^{\circ}$ with $\theta_{13}/\bar{\theta}_{13}$ priors. Lower left panel: same as the upper right one but without $\theta_{13}/\bar{\theta}_{13}$ priors. Lower right panel: same as the upper right one but only $\bar{\theta}_{13}$ prior is included.
@@ 4.2 Results for the CPT framework
We now present our numerical results. In Fig. 7 we show the results for $x-\bar{x}$ at 90% C.L. for $x=\delta_{CP},~{}\theta_{13},~{}\theta_{23}$ and both mass splittings assuming $\delta_{CP}^{\text{true}}=0,~{}-90^{\circ}$ and no CPT violation. Since a LBL experiment is not very sensitive to the solar mixing angle $\theta_{12}$, we do not show the CPT results for it for simplicity. In general, the addition of a NF to DUNE improves the precision on the oscillation parameters for neutrinos and anti-neutrinos simultaneously, in particular when CID is implemented. An exception to this is the results for $\Delta m_{31}^{2}$ which are similar for 20 years of DUNE and DUNE+NF, nearly independent of the choice of CID. We find that the neutrino oscillation parameters can be more precisely constrained than the anti-neutrino parameters due to the larger statistics.
@@
Accelerator experiments do not have a large sensitivity to the solar parameters, they are most precisely determined by solar and reactor experiments. Concerning CPT violation we see that DUNE alone cannot exclude non-zero $\bar{\Delta}m_{21}^{2}$ at 90% C.L. while this parameter in neutrino mode can be constrained to be non-zero. Adding a neutrino factory, in particular with the FNAL baseline allows to exclude zero $\bar{\Delta}m_{21}^{2}$.
@@
For $\theta_{13}$ DUNE only cannot distinguish $\theta_{13}=8.5^{\circ}$ from $\theta_{13}=9.5^{\circ}$ at 90% C.L. under the assumption of CPT violation even when doubling the running time. This is because the free $\theta_{23}$ angle and CP phase can degenerate with $\theta_{13}$. As shown in the left panel of Fig. 9, for $\nu_{\mu}\rightarrow\nu_{e}$ and $\nu_{\mu}\rightarrow\nu_{\mu}$ channels which are main signals at DUNE, the oscillation probabilities with $(\theta_{13}=9.5^{\circ}$, $\theta_{23}=43^{\circ}$, $\delta_{CP}=-83^{\circ})$ are very close to that of $(\theta_{13}=8.5^{\circ},\theta_{23}=49^{\circ},\delta_{CP}=-90^{\circ})$, leading to the island-like structure in Fig. 7. However, the $\nu_{e}\rightarrow\nu_{e}$ channel is almost independent of the CP phase and $\theta_{23}$ angle, which means a measurement of $\nu_{e}$ disappearance channel can significantly improve the $\theta_{13}$ sensitivity, in a similar fashion to a typical medium baseline reactor neutrino experiment, except with the added channel of $\nu_{e}\to\nu_{e}$ instead of just $\bar{\nu}_{e}\to\bar{\nu}_{e}$. The precision on $\bar{\theta}_{13}$ is ±$\pm 2.5^{\circ}$. The addition of a NF excludes the $\theta_{13}=9.5^{\circ}$ and the precision on $\bar{\theta}_{13}$ also improves.
@@
For $\bar{\theta}_{23}$ DUNE alone or combined with a NF cannot resolve the octant at 90% C.L. in the CPT violating scenario unless the NF has good $\mu$CID. As shown in the right panel of Fig. 9, the $\nu_{\mu}$ disappearance channel is sensitive to $\sin^{2}2\theta_{23}$ and hence not sensitive to the $\theta_{23}$ octant Hyper-Kamiokande:2018ofw . However, $\nu_{\mu}$ appearance is sensitive to $\sin^{2}\theta_{23}$, so the octant can be resolved. In the case of neutrino factory, $\mu$CID is required to provide a clean measurement of the $\nu_{\mu}$ appearance channel. Consequently, the sensitivity of $\theta_{23}$ octant can be largely improved. The octant can be resolved for the neutrino parameter due to the larger statistics in the neutrino channel.
@@
Of interest is also the impact of $\theta_{13}$ in the determination of $\theta_{23}$ octant because the main contribution to the oscillation probability $P_{\mu e}$ is $4\sin^{2}\theta_{23}\sin^{2}\theta_{13}\cos^{2}\theta_{13}\sin^{2}\Delta_{31}$ with $\Delta_{31}\equiv\Delta m^{2}_{31}L/4E_{\nu}$ Hyper-Kamiokande:2018ofw , which means $\theta_{13}$ can affect the $\theta_{23}$ octant. In Fig. 8 we study the expected constraints without using a prior on neither $\theta_{13}$ nor $\bar{\theta}_{13}$. DUNE alone cannot resolve the octant for neither $\theta_{23}$ nor $\bar{\theta}_{23}$. When a NF is added only two degenerate solutions for $\bar{\theta}_{23}$ survive due to the less statistics in the antineutrino mode. The prior on $\bar{\theta}_{13}$ is the driving force behind removing the wrong octant solutions for $\theta_{23}$. Having a prior on $\bar{\theta}_{13}$ also allows to remove the wrong-octant solution for $\bar{\theta}_{23}$ for a NF with $\mu$CID as discussed above.
@@
So far there is no experimental evidence that CPT is violated in nature. However, there is a mild discrepancy in the measurement of $\theta_{13}$ and $\bar{\theta}_{13}$ coming from the measurement of solar Super-Kamiokande:2023jbt and accelerator T2K:2023smv neutrinos compared to reactor antineutrinos DayaBay:2022orm ; Shin:2020mue ; Soldin:2024fgt . We now consider benchmark values of $\sin^{2}\theta_{13}=0.032,~{}\sin^{2}\bar{\theta}_{13}=0.022$, which are consistent with the solar neutrino measurement and reactor neutrino measurement, respectively. To study the impact of CPT for $\theta_{13}$ we show in Fig. 10 the expected sensitivities to $\theta_{13}$ and $\bar{\theta}_{13}$ assuming the true values to be CPT violating. We again see that the addition of a NF excludes the higher solution for $\theta_{13}$ and the precision on both parameters is drastically improved due to the measurement of $\nu_{e}$ disappearance channel as discussed above; this would allow for the discovery of CPT violation if this hint was realized in nature. We now consider a similar scenario but with the octant of $\theta_{23}$. If we assume that $\theta_{23}^{\text{true}}=49^{\circ},~{}\bar{\theta}_{23}^{\text{true}}=42^{\circ}$ we again find that having a prior on $\bar{\theta}_{13}$ is crucial to exclude the wrong-octant solutions with a NF. In some cases a CPT violating signal in the octant could be discovered with a NF with given the inclusion of other oscillation data.

### v2

@@ 3.2 Results for non-standard vector interactions
For a more general case, we take the off-diagonal NSI parameters to be complex and marginalize over the corresponding phases in Fig. 3. Due to the free complex phase, the degeneracies can be enlarged and significantly reduce the experimental sensitivities to the NSI parameters. Notably, the degeneracy between $\epsilon_{ee}-\epsilon_{\mu\mu}$ and $|\epsilon_{e\tau}|$ leads to a wing-like structure for the original DUNE configuration which comes from the non-trivial dependence among the $\phi_{e\tau}$, $\delta_{CP}$ and $\theta_{23}$ parameters Coloma:2015kiu . To better understand this degeneracy we show in the left panel of Fig. 4 the oscillation probabilities without NSI and with non-zero $\epsilon_{ee}-\epsilon_{\mu\mu}$ and $\epsilon_{e\tau}$. For the standard case without NSI, we use the values of oscillation parameters from Tab. 1, while we take the best fit values that minimize the $\chi^{2}$ function for the NSI case. More explicitly, we choose the point ($\epsilon_{ee}-\epsilon_{\mu\mu}=1.2$ and $|\epsilon_{e\tau}|=0.3$) in the wing-like structure panel of Fig. 3 as the input. The $\chi^{2}$ minimum is achieved for the parameter $\phi_{e\tau}=-152^{\circ}$, $\theta_{23}=42^{\circ}$, and $\delta_{CP}=-53^{\circ}$, which indicates this parameter set has a degeneracy with the standard case without NSI. As DUNE only measures $\nu_{e}$ appearance and $\nu_{\mu}$ disappearance channels (and their anti-neutrino equivalents), this degeneracy significantly reduces DUNE’s sensitivity to these NSI parameters. The addition of a neutrino factory provides powerful new measurements to break this degeneracy and significantly enhance the sensitivity in the $\epsilon_{ee}-\epsilon_{\mu\mu}$-$|\epsilon_{e\tau}|$ parameter space beyond the naive expectations from statistics and new oscillation channels alone. This is because a neutrino factory can measure the $\nu_{e}$ disappearance channel and the $\nu_{\mu}$ appearance channel, which help break the existing degeneracy in the $\nu_{e}$ appearance and $\nu_{\mu}$ disappearance channels.
@@
Figure 3: The experimental constraint on different combinations of two NSI parameters at 90% confidence level, where the corresponding complex phase of NSI off-diagonal parameter and $\delta_{\rm CP}$ are marginalized over. See Fig. 13 in the appendix for the sensitivities for the BNL configuration.
@@
In addition to the expected constraints on the NSI parameters, we also study the precision of the Dirac CP phase as well as the sensitivity to $\theta_{23}$ in the presence of NSI. To quantify the expected precision on to $\delta_{CP}$, we define $\Delta\delta_{CP}\equiv\rm max\left\{|\delta_{CP}^{\rm fit}-\delta_{CP}^{\rm true}|\right\}$ to be the maximum absolute value of the difference between $\delta_{CP}^{\rm fit}$ and $\delta_{CP}^{\rm true}$, where $\delta_{CP}^{\rm fit}$ is the fitted Dirac CP phase which gives $\Delta\chi^{2}=2.71$ (90% C.L.) and $\delta_{CP}^{\rm true}$ is the fixed true Dirac CP phase. We take the benchmark value of $\delta^{\rm true}_{CP}=0^{\circ}$ or $-90^{\circ}$, and three benchmark values of $\theta^{\rm true}_{23}=42^{\circ}$, $45^{\circ}$, or $49^{\circ}$ for completeness. Similar to the standard oscillation framework without NSI, the precision on $\delta_{CP}^{\rm true}=0^{\circ}$ is better than for the $\delta_{CP}^{\rm true}=-90^{\circ}$ case while the octant of $\theta_{23}$ has only a mild impact on the CP precision. A sizable improvement exists for $\delta_{CP}^{\rm true}=-90^{\circ}$, as seen in Fig. 5. The results for $\delta^{\rm{true}}_{CP}=0$ are shown in the appendix which show qualitatively similar improvements to the precision of $\delta_{CP}$ with the addition of a neutrino factory.
@@
In Fig. 6 we show the sensitivity to $\theta_{23}$ for different benchmark values of $\delta^{\rm true}_{CP}$ and $\theta^{\rm true}_{23}$ as discussed above. Here we take the $\epsilon_{e\mu}$ parameter for illustration since the neutrino factory has a significant improvement in this case, while the results for other NSI parameters are shown in the appendix. For this the figure, we assume $\epsilon_{e\mu}=0$. We find that the sensitivity to $\theta_{23}$, for the benchmark values of $\theta_{23}^{\rm true}=42^{\circ}$ and $49^{\circ}$, at DUNE can be significantly reduced if one also considers NSI. As a result, the octant of $\theta_{23}$ cannot be determined with over $\sim 3\sigma$ C.L. at the benchmarks considered. With the addition of a neutrino factory, however, the sensitivity to the octant of $\theta_{23}$ can reach over $5\sigma$. We also find that the true value of $\delta_{CP}^{\text{true}}$ has only a small impact on the sensitivity.
@@
Figure 5: The expected uncertainty on $\delta_{CP}$ at 90% confidence level, where the corresponding NSI parameter is minimized over. In this figure, the off-diagonal NSI parameters are taken as complex and the corresponding phase is minimized together with its modulus at a time. We assume $\delta_{CP}^{\text{true}}=-90^{\circ}$ and $\theta_{23}^{\text{true}}=42^{\circ},~45^{\circ},~49^{\circ}$. See Fig. 14 in the appendix for the BNL configuration and Fig. 18 in the appendix for $\delta_{CP}=0$.
@@
Figure 6: The experimental sensitivity to $\theta_{23}$ for different true values of $\theta_{23}$ and we assume no NSI effect in the true case. For the fitting parameter, we consider the complex off-diagonal NSI parameter $\epsilon_{e\mu}$, which is minimized over including its complex phase. See Figs. 19-23 in the appendix for the BNL configuration and other NSI parameter setups.
@@ 4.1 Framework
The CPT symmetry, which connects the three discrete symmetries: charge conjugation (C), parity (P), and time reversal (T), is one of the most essential symmetries in particle physics and its invariance has been a guiding tool to build models. The CPT theorem is based on several basic assumptions on models, including Lorentz invariance, hermiticity of the Hamiltonian, and local commutativity. Despite its foundational nature, CPT can be violated in several classes of more complete models, some of which affect neutrino oscillations Kostelecky:1988zi ; Kostelecky:1989jp ; Kostelecky:1989jw ; Sheikh-Jabbari:2000hdt ; Carroll:2001ws ; Greenberg:2002uu ; Barenboim:2002tz ; DeGouvea:2002xp ; Ge:2019tdi . Certainly, the discovery of the CPT violation would indicate that at least one basic assumption should be rejected and all our model building strategy need to be revisited. Therefore, testing the CPT symmetry is one of the major tasks in particle physics.
@@
Here $x$ ($\bar{x}$) corresponds to the oscillation parameter set in the neutrino (anti-neutrino) oscillation probabilities.
Figure 7: The experimental sensitivity to the difference between the oscillation parameters for neutrino and anti neutrinos $x$-$\bar{x}$ at 90% confidence level for different experimental setups assuming CPT conservation. We mark the best-fit value with a star. The dashed line shows the CPT conserving line. See Fig. 15 in the appendix for the BNL configuration.
@@
Figure 9: The expected experimental constraint on $\theta_{23}$-$\bar{\theta}_{23}$ at 90% confidence level. Left panel: both $\theta_{13}$ and $\bar{\theta}_{13}$ are free parameters without any prior. Right panel: both $\theta_{13}$ and $\bar{\theta}_{13}$ are free parameters, but only the prior of $\bar{\theta}_{13}$ is included. See Fig. 16 in the appendix for the BNL configuration.
Figure 10: CPT violation benchmark plot. Upper left panel: $\sin^{2}\theta_{13}^{\rm true}=0.032,\sin^{2}\bar{\theta}_{13}^{\rm true}=0.022$. Upper right panel: $\theta_{23}^{\rm true}=49^{\circ},\bar{\theta}_{23}^{\rm true}=42^{\circ}$ with $\theta_{13}/\bar{\theta}_{13}$ priors. Lower left panel: same as the upper right one but without $\theta_{13}/\bar{\theta}_{13}$ priors. Lower right panel: same as the upper right one but only $\bar{\theta}_{13}$ prior is included. See Fig. 17 in the appendix for the BNL configuration.
@@ 4.2 Results for the CPT framework
We now present our numerical results on CPT violation sensitivities. In Fig. 7 we show the results for $x-\bar{x}$ at 90% C.L. for $x=\delta_{CP},~\theta_{13},~\theta_{23}$ and both mass splittings assuming $\delta_{CP}^{\text{true}}=0$ or $-90^{\circ}$ and no CPT violation. Since a long-baseline experiment is not very sensitive to the solar mixing angle $\theta_{12}$, we do not show the CPT results for it for simplicity. In general, the addition of a NF to DUNE improves the precision on the oscillation parameters for neutrinos and anti-neutrinos simultaneously, in particular when CID is implemented. An exception to this is the results for $\Delta m_{31}^{2}$ which are similar for 20 years of DUNE and DUNE+NF, nearly independent of the choice of CID. We find that the neutrino oscillation parameters can be more precisely constrained than the anti-neutrino parameters due to the larger statistics.
@@
Accelerator experiments do not have a large sensitivity to the solar parameters; they are most precisely determined by solar and reactor experiments. Concerning CPT violation we see that DUNE alone cannot exclude non-zero $\bar{\Delta}m_{21}^{2}$ at 90% C.L. while this parameter in neutrino mode can be constrained to be non-zero. Adding a neutrino factory, in particular with the FNAL baseline allows to exclude zero $\bar{\Delta}m_{21}^{2}$.
@@
For $\theta_{13}$ DUNE only cannot distinguish $\theta_{13}=8.5^{\circ}$ from the $\theta_{13}=9.5^{\circ}$ degenerate point at 90% C.L. under the assumption of CPT violation even when doubling the running time. This is because the free $\theta_{23}$ angle and CP phase can degenerate with $\theta_{13}$. As shown in the left panel of Fig. 8, for $\nu_{\mu}\rightarrow\nu_{e}$ and $\nu_{\mu}\rightarrow\nu_{\mu}$ channels which are main signals at DUNE, the oscillation probabilities with $(\theta_{13}=9.5^{\circ}$, $\theta_{23}=43^{\circ}$, $\delta_{CP}=-83^{\circ})$ are very close to that of $(\theta_{13}=8.5^{\circ},\theta_{23}=49^{\circ},\delta_{CP}=-90^{\circ})$, leading to the island-like structure in Fig. 7. However, the $\nu_{e}\rightarrow\nu_{e}$ channel is almost independent of the CP phase and $\theta_{23}$ angle, which means a measurement of $\nu_{e}$ disappearance channel can significantly improve the $\theta_{13}$ sensitivity, in a similar fashion to a typical medium baseline reactor neutrino experiment, except with the added channel of $\nu_{e}\to\nu_{e}$ instead of just $\bar{\nu}_{e}\to\bar{\nu}_{e}$. The precision on $\bar{\theta}_{13}$ is ±$\pm 2.5^{\circ}$. The addition of a NF excludes the $\theta_{13}=9.5^{\circ}$ degeneracy and the precision on $\bar{\theta}_{13}$ also improves.
@@
For $\bar{\theta}_{23}$ DUNE alone or combined with a NF cannot resolve the octant at 90% C.L. in the CPT violating scenario unless the NF has good $\mu$CID. As shown in the right panel of Fig. 8, the $\nu_{\mu}$ disappearance channel is sensitive to $\sin^{2}2\theta_{23}$ and hence not sensitive to the $\theta_{23}$ octant. However, appearance channels are sensitive to $\sin^{2}\theta_{23}$, so the octant can be resolved. In the case of a neutrino factory, $\mu$CID is required to provide a clean measurement of the $\nu_{\mu}$ appearance channel. Consequently, the sensitivity of $\theta_{23}$ octant can be largely improved in the presence of CPT violation. The octant can be resolved for the neutrino parameter due to the larger statistics in the neutrino channel.
@@
Of interest is also the impact of $\theta_{13}$ in the determination of $\theta_{23}$ octant because the main contribution to the oscillation probability $P_{\mu e}$ is $4\sin^{2}\theta_{23}\sin^{2}\theta_{13}\cos^{2}\theta_{13}\sin^{2}\Delta_{31}$ with $\Delta_{31}\equiv\Delta m^{2}_{31}L/4E_{\nu}$, which means $\theta_{13}$ can affect the $\theta_{23}$ octant. In Fig. 9 we study the expected constraints without using an external prior on $\theta_{13}$ or $\bar{\theta}_{13}$. DUNE alone cannot resolve the octant for either $\theta_{23}$ or $\bar{\theta}_{23}$. When a NF is added only two degenerate solutions for $\bar{\theta}_{23}$ survive due to the lower statistics in the anti-neutrino channel. The prior on $\bar{\theta}_{13}$ from reactor anti-neutrino measurements is the driving force behind removing the wrong octant solutions for $\theta_{23}$. Having a prior on $\bar{\theta}_{13}$ also allows to remove the wrong-octant solution for $\bar{\theta}_{23}$ for a NF with $\mu$CID as discussed above.
@@
So far there is no experimental evidence that CPT is violated in nature. However, there is a mild discrepancy in the measurement of $\theta_{13}$ and $\bar{\theta}_{13}$ coming from the measurement of solar Super-Kamiokande:2023jbt and accelerator T2K:2023smv neutrinos compared to reactor anti-neutrinos DayaBay:2022orm ; Shin:2020mue ; Soldin:2024fgt . We now consider benchmark values of $\sin^{2}\theta_{13}=0.032,~\sin^{2}\bar{\theta}_{13}=0.022$, which are consistent with the solar neutrino measurement and reactor neutrino measurement, respectively. To study the impact of CPT for $\theta_{13}$ we show in Fig. 10 the expected sensitivities to $\theta_{13}$ and $\bar{\theta}_{13}$ assuming the true values to be CPT violating. We again see that the addition of a NF excludes the higher solution for $\theta_{13}$ and the precision on both parameters is drastically improved due to the measurement of $\nu_{e}$ disappearance channel as discussed above; this would allow for the discovery of CPT violation if this hint was realized in nature. We now consider a similar scenario but with the octant of $\theta_{23}$. If we assume that $\theta_{23}^{\text{true}}=49^{\circ},~\bar{\theta}_{23}^{\text{true}}=42^{\circ}$ we again find that having a prior on $\bar{\theta}_{13}$ is crucial to exclude the wrong-octant solutions with a NF. In some cases a CPT violating signal in the octant could be discovered with a NF given the inclusion of other oscillation data.
Finally, we comment on existing constraints from oscillations. Long-baseline experiments will not be competitive with the measurements of the solar parameters $\theta_{12}$ and $\Delta m^{2}_{21}$ relative to the combined analysis of solar and reactor data which have more precision in the standard case Denton:2023zwa and have the benefit of one class of experiments measuring neutrinos (solar) and the other measuring antineutrinos (reactors) Barenboim:2023krl . DUNE and a NF will improve on the constraints for the other four parameters considerably as shown in fig. 7. There are no constraints now on CPT violation appearing in $\delta$ as $\delta$ has not yet been measured in the CPT conserving case, so the sensitivity of LBL experiments is world-leading. As for the other three parameters, the existing constraints at 90% are at $|\Delta\sin^{2}\theta_{23}|\lesssim 0.13$, $|\Delta\sin^{2}\theta_{13}|\lesssim 0.019$, and $|\Delta(\Delta m^{2}_{31})|\lesssim 0.14\times 10^{-3}$ eV2 Tortola:2020ncu where $\Delta x\equiv x-\bar{x}$. In comparison, we find that long-baseline accelerator experiments can achieve $|\Delta\sin^{2}\theta_{23}|\lesssim 0.02$, $|\Delta\sin^{2}\theta_{13}|\lesssim 0.007$, and $|\Delta(\Delta m^{2}_{31})|\lesssim 0.03\times 10^{-3}$ eV2 where we have assumed that the relevant degeneracies can be lifted via the inclusion of data for $\bar{\theta}_{13}$.